16. What is the concentration (M) of KCl in a solution made by mixing 25.0mlL ofo,100 at is the concentration (M) of KCI in a solution made by mixing 25.0 mL of 0.100 M KCI with 500 mL of 0.100 MKCI? 0.00」Thol A) 0.100 B) 0.0500 C) 0.0333 D) 0.0250 E) 125
Answer #1

Answer:

Given first solution, 25 ml of 0.1 M of KCl, moles of KCl=(0.1 mol/L)*(0.025 L)=0.0025 mol

Second solution, 50 ml of 0.1 M of KCl, moles of KCl=(0.1 mol/L)*(0.050 L)=0.0050 mol

If we add two solution, third solution formed. Number of moles of third solution=0.0025 mol+0.0050 mol=0.0075 mol and total volume V=0.025 L+0.050L=0.075 L

Molarity of third solution=0.0075 mol/0.075 L=0.1 M